OpenStudy (anonymous):

A police car is stopped at a set of lights. the bat mobile speeds passed the police car at 100 km/h in a school zone. If the police car can accelerate at 3.6 m/s^2 , how long does it take the police car to catch up to the batmobile?

OpenStudy (anonymous):

how do i do it

OpenStudy (anonymous):

it's physics, sorry i put it in the wrong place, but whatever whoever can answer this i'll give a medal to anyways

OpenStudy (anonymous):

grade 11 physics, that is

OpenStudy (anonymous):

The distance traveled by both after 't' is going to be equal. So equate the two distance equations to find 't'. I hope you can figure out which formulas to use :)

OpenStudy (anonymous):

um

OpenStudy (anonymous):

ok so i will make both d equal

OpenStudy (anonymous):

you'll use two distance equations and equate them

OpenStudy (anonymous):

so initial velocity is 0 m/s right? what is final velocity? 27.8 m/s ?

OpenStudy (anonymous):

which equations of motions are you planning to use?

OpenStudy (anonymous):

final velocity squared = initial V squared + 2ad

OpenStudy (anonymous):

um idk about other one yet

OpenStudy (anonymous):

what should i use?

OpenStudy (anonymous):

distance moved by the bat mobile=d1=velocity*time and distance moved by the police car=d2=vi*t + 1/2*a*t^2 since d1=d2... put in value and solve to find the value of 't'

OpenStudy (anonymous):

but we dont have v2 ... do we? ???

OpenStudy (anonymous):

we dont have t either

OpenStudy (anonymous):

\[v_f^2 = v_i^2 + 2ad\]

OpenStudy (anonymous):

we can rearrange it and get \[vf^2 / 2a since v_i is = 0\]

OpenStudy (anonymous):

we don't need v2 and we have to find t

OpenStudy (anonymous):

could you solve the equation in front of me.?

OpenStudy (anonymous):

d1=d2 27.8*t=(0)+1/2(3.6)t^2

OpenStudy (anonymous):

what is that first thing you put, the 27.8? is it v2 or v1?

OpenStudy (anonymous):

100km/hr converted to m/s

OpenStudy (anonymous):

i know but is that the variable of final velocity or initial velocity?!?!

OpenStudy (anonymous):

i have an exam tomorrow end of semester exam

OpenStudy (anonymous):

that is the velocity of the batmobile

OpenStudy (anonymous):

i know but are we replacing it for the variable V1 OR V2?

OpenStudy (anonymous):

we have two variables of velocity. which one is 27.8?

OpenStudy (anonymous):

?!?!

OpenStudy (anonymous):

In this question we don't' know that final and initial velocity of the bat mobile so we just presume it as the constant velocity.and so we use the basic formula of velocity*time=distance